With following values, c = 0,7, t = 0,41 and m = 0,86 use the multiplier and calculate the potential impact of the President’s R791 billion infrastructure investment plan on the value of GDP in 2021. 


1
Expert's answer
2021-04-26T21:12:28-0400

the formula of multipliers ="\frac{dy}{d}=[\frac{1}{1-b(1-t)+m}]"


let's say investment increases by 791

then

"\frac{dy}{791}=[\frac{1}{1-0.7(1-0.41)+0.86}]"


"\frac{dy}{791}=[\frac{1}{1-0.7(0.59)+0.86}]"


"\frac{dy}{791}=[\frac{1}{1-0.413+0.86}]"


"\frac{dy}{791}=[\frac{1}{1.447}]"


"\frac{dy}{791}=0.691"


"dy=546.648"


GDP in 2021 increases by 546.648 bn